Consider the reaction 

8H₂S(g)+4O₂(g)→8H₂O(g)+S₈(g) 

Δ[H₂S]/Δt = -0.067 M/s 

You may want to reference (Pages 587 - 592) Section 14.3 while completing this problem. 


Part A 

Find Δ[O₂]/Δt. 

Express your answer to two significant figures and include the appropriate units. 


Part B 

Find Δ[H₂O]/Δt. 

Express your answer to two significant figures and include the appropriate units. 
Answer #1

rate of reaction is given as:
rate = sign * (1/x) * Δ[concentration]/Δt

sign is negative for reactant and positive for product
x is stoichiometric coefficient in chemical equation


In terms of H2S:
rate = (-1/8) * Δ[H2S]/Δt
rate = (-1/8) * (-0.067 M/s)
= 8.375*10^-3 M/s

a)
In terms of O2,
Rate = (-1/4) * Δ[O2]/Δt
8.375*10^-3 M/s = (-1/4)*Δ[O2]/Δt
Δ[O2]/Δt = 0.0335 M/s
Answer: 0.034 M/s

b)
In terms of H2O:
rate = (-1/8) * Δ[H2O]/Δt
8.375*10^-3 M/s = (-1/8) * Δ[H2O]/Δt
Δ[H2O]/Δt = 0.067 M/s
Answer: 0.067 M/s